Find values for m and b in the following system so that the solution to the system is (-3,4).
5x+7y=b
mx+y=22

Answer by stanbon(75887)

Find values for m and b in the following system so that the solution to the system is (-3,4).
5x+7y=b
mx+y=22
------------
Substitute the point values into the equations:
5*-3 + 7*4 = b
b = -15 + 28 = 13
-------------
-3m + 4 = 22
-3m = 18
m = -6
---------------
Final solution: b=13 ; m=-6
================
Cheers,
Stan H.